OpenStudy (imamermaid29):

can someone please help me? Find the value for x that makes the equation 1.5x + 6 = 3(2.5x - 6) true.

OpenStudy (misty1212):

HI!! \[ 1.5x + 6 = 3(2.5x - 6) \]

OpenStudy (misty1212):

multiply out, get \[1.5x+6=7.5x-18\]

OpenStudy (misty1212):

subract \(1.5x\)from both sides, get \[6=6x-18\]

OpenStudy (misty1212):

then solve in two steps a) add 18 b) divide by 6

OpenStudy (imamermaid29):

THanks :) @misty1212

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]